100 liters of a 50% solution is obtained by mixing a 60% solution with a 20% solution. How many liters of each solution must be used to obtain the desired mixture?
Use linear equations please.. :)

Answer by fractalier(6550)

Call the amount of 60% solution you use, x.
Then the amount of 20% solution is 100-x.
The set up looks like this
.60x + .20(100-x) = .50(100)
.60x + 20 - .20x = 50
.40x = 30
x = 75 liters at 60% and then
100-x = 25 liters at 20%